OpenStudy (anonymous):

Bradley has 3 square pieces of cardboard with each side equal to x units. For each piece, he does something different to it according to each part below: Part A: Bradley pasted rectangular strips along two adjacent sides of the cardboard to increase its length and width by y units each. What will be the change in the area of the piece of cardboard? Show your work. (4 points) Part B: Bradley cut off rectangular strips from two adjacent sides of the cardboard to decrease its length and width by y units each. What will be the change in the area of the piece of cardboard? Show your work.

OpenStudy (anonymous):

Part C: Bradley cut off a strip from one side of the cardboard and pasted the strip on an adjacent side of the cardboard to increase its length by y units and decrease its width by y units. What will be the change in the area of the cardboard? Show your work

OpenStudy (anonymous):

@perl @satellite73 @shrutipande9 @dan815

OpenStudy (anonymous):

1.So, the dimensions of the square are x and x Then,the area of the square piece=x*x=x^2 a)New length =x+y New breadth=x+y New area=(x+y)(x+y)=(x+y)^2 =x^2+2xy+y^2 { as (a+b)=a^2+b^2+2ab} Change in area=new area-old area =x^2+2xy+y^2-x^2 =2xy+y^2 b)New length =x-y new breadth=x-y new area=(x-y)(x-y)=(x-y)^2 {as (a-b)^2=a^2+b^2-2ab} =x^2-2xy+y^2 Change in area=new area-old area =x^2-2xy+y^2-x^2 =y^2-2xy c)New lenght=x+y new breadth=x-y new area=(x+y)(x-y) {as (a+b)(a-b)=a^2-b^2} =x^2-y^2 change in area=x^2-y^2-x^2 =-y^2 2.Length of rectangle = (4*2+2x-1) =8-1+2x =7+2x Width of the rectangle=3*3-x+4 =9+4-x =13-x a)area of the rectangle=l*b =(7+2x)(13-x) =7(13-x)+2x(13-x) =91-7x+26x-2x^2 =91+19x-2x^2 unit^2 b)Yes, as (7+2x)(13-x)=(13-x)(7+2x)=91+19x-2x^2 ( you can do the calculations if you want to check it.) c)The degree of polynomial is 2 as the highest power is 2 and thus it is a quadratic polynomial (based upon its degree).You can also classify it as trinomial based upon the number of terms it contains. Hope this helps and u understand all the identities i have used here.Thank U.

OpenStudy (anonymous):

The area of each original square is : A = x^2 ........... A = s^2........... s= side of the square A) L = x + y and W= x + y Area = ( x + y )^2 = x^2 + 2xy + y^2 Increase in area : Ad= x^2 + 2xy + y^2 - x^2 = 2xy + y^2 B ) The figure is a square of side ( x - y ) its area is A = ( x - y )^2 = x^2 - 2xy + y^2 The change in area is : Ad = x^2 - 2xy + y^2 - x^2 = y^2 - 2xy c) The rectangle is L = x + y and W= x - y The area of the rectangle is A = L x W Then the area is A = ( x + y )( x - y ) = x^2 - y^2 The new area is; x^2 - y^2 - x^2 = - y^2 .......... Negative means a decrease The new area decreases by y^2 2) A) The area is A= L x W ... A = ( 4x^2 + 2x - 1 )( 3x^3 -x + 4 ) = 12x^5 - 4x^3 + 16x^2 + 6x^4 -2x^2 + 8x - 3x^3 +x - 4 ... A = 12x^5 + 6x^4 -7x^3 + 14x^2 + 9x - 4........... AREA B) Yes, polynomials are closed over multiplication because the product of two polynomials is a polynomial. C) It is a polynomial of 5th degree in one variable.

OpenStudy (anonymous):

i dont undeerstand your expplantion

OpenStudy (anonymous):

@perl

OpenStudy (perl):

lets start with a drawing

OpenStudy (perl):



OpenStudy (perl):